# Replay-Free Continual Low-Rank Adaptation with Dynamic Memory

### v3

@@ Abstract
We revisit continual learning (CL), which enables pre-trained vision transformers (ViTs) to sequentially fine-tune on new downstream tasks over time. However, as the scale of these models increases, catastrophic forgetting remains a more serious challenge. Recent studies highlight a crossover between CL techniques and parameter-efficient fine-tuning (PEFT), which focuses on fine-tuning only a small set of trainable parameters to adapt to downstream tasks, such as low-rank adaptation (LoRA). While LoRA achieves faster convergence and requires fewer trainable parameters, it has seldom been explored in the context of continual learning. To address this gap, we propose a novel PEFT-CL method called Dual Low-Rank Adaptation (DualLoRA), which introduces both an orthogonal LoRA adapter and a residual LoRA adapter parallel to pre-trained weights in each layer. These components are orchestrated by a dynamic memory mechanism to strike a balance between stability and plasticity. Additionally, we propose a scheme to predict task identity with confidence and calibrate the model’s outputs accordingly. On ViT-based models, we demonstrate that DualLoRA offers significant advantages in accuracy, inference speed, and computation efficiency in training over existing CL methods across multiple benchmarks.
@@ 1 Introduction
Continual learning (CL) [22], which involves training a model on a sequence of tasks, often faces the challenge of catastrophic forgetting—a significant decline in performance on previously learned tasks when learning new ones. This issue persists even in the continual training of vision foundation models [25, 26], despite their strong generalization capabilities and robustness. Existing replay-based CL schemes [1] typically mitigate forgetting by storing a subset of data from previous tasks as exemplars. However, this approach can be impractical due to storage constraints and data retention policies. Meanwhile, architecture-based CL schemes [14] allocate isolated parameters for each task, using a task identifier during training to prevent interference across tasks. Nevertheless, modifying the architecture of vision foundation models for fine-tuning on downstream tasks can compromise their pre-trained capabilities, as the pre-trained weights are learned within a fixed architecture. Furthermore, these methods often assume that task identity is known during inference, which is unrealistic in real-world scenarios.
@@
Figure 1: PEFT-based continual leanring schemes dominate the ImageNet-R dataset in recent years.
@@
Recently, parameter-efficient fine-tuning (PEFT) techniques have gained significant attention for their ability to adapt foundation models to downstream tasks by updating or adding only a small number of parameters. Moreover, PEFT methods have demonstrated notable robustness in mitigating catastrophic forgetting during fine-tuning on sequential tasks. In particular, prompt-tuning [25, 26, 20] and low-rank adaptation (LoRA) [12] are two widely used PEFT approaches that have achieved remarkable success in continual learning, significantly outperforming conventional CL schemes across established benchmarks. Prompt-tuning focuses on learning a prompt pool, which enables the matching of an image with a set of prompt vectors and the alignment of image features with patch tokens. However, existing prompt-based CL methods require using the original pre-trained encoder as the query function. Consequently, image tokens must be processed through the network twice, resulting in substantial computational overhead and increased inference latency.
In contrast, LoRA methods require fewer trainable parameters to achieve comparable performance on domain-specific tasks and enable faster inference speeds than prompt-tuning. Nevertheless, vanilla LoRA [7] struggles to preserve satisfactory performance in continual learning settings due to significant interference across different tasks. Inspired by gradient projection techniques [19], InfLoRA [12] takes the first step in mitigating such interference by initializing LoRA adapter parameters for new tasks within a subspace orthogonal to the gradient subspace of previously learned tasks. However, InfLoRA requires passing through the encoder twice for each sample in a task: the first pass extracts the gradient subspace for LoRA initialization, while the second pass updates the parameters, leading to substantial computational overhead in the training phase.
To this end, we propose a novel continual learning method, dual low-rank adaptation (DualLoRA), which integrates an orthogonal adapter and a residual adapter into each layer of pre-trained vision transformers (ViTs). Specifically, the orthogonal adapter $\mathbf{O}$ is updated exclusively in directions orthogonal to features extracted from previously learned tasks, while the residual adapter $\mathbf{R}$ is updated in a task-specific subspace spanned by the residual bases extracted from previously learned tasks. This design enhances stability, i.e., robustness against forgetting in old tasks, through orthogonal adapters while improving plasticity, i.e., the ability to adapt continuously to new tasks, through residual adapters—thus achieving a balance between both objectives. Unlike InfLoRA, DualLoRA efficiently extracts orthogonal bases from the feature subspace of previously learned tasks and projects the updates of $\mathbf{O}$ and $\mathbf{R}$ using matrices constructed from these extracted bases. Moreover, these bases can be used to compute task relevance during inference, allowing the residual adapter’s outputs to be adjusted based on task relevance, thereby mitigating components that could degrade test performance. This mechanism, which dynamically modulates the parameters of the residual adapter based on input test samples during inference, is referred to as dynamic memory (DM), as illustrated in Fig. 2.
Extensive experimental results demonstrate that DualLoRA outperforms existing PEFT methods across various continual learning benchmarks, without incurring significant additional computational or memory overhead. The main contributions of this paper are summarized as follows:
@@ 2.1 Gradient Projection in CL
Gradient projection is widely employed in continual learning to mitigate catastrophic forgetting by updating parameters in directions that minimize interference with previously learned tasks. OGD [3] was the first to implement gradient descent in the direction orthogonal to the stored gradient directions computed in previous tasks. The follow-up work GPM [19] extracts orthogonal bases of task representations from randomly selected training data via singular value decomposition (SVD). A subsequent study, TRGP [13], introduces the concept of a trust region, allowing partial reuse of selected bases from previous tasks. FSDGPM [2] further evaluates the importance of bases in GPM by assessing the sharpness of the loss landscape, assigning weights to the bases in the projection matrix according to their relative importance. Due to the computational expense of determining loss landscape sharpness, SGP [18] offers an alternative by using accumulated singular values as an importance indicator to scale the projection matrix in GPM. Additionally, several studies [23, 9, 11, 27] have explored relaxing the orthogonality constraints and optimizing the relaxation factor to enhance performance. However, these approaches are primarily developed for simpler models such as CNNs and face significant challenges when applied to advanced architectures like ViTs. For instance, the high dimensionality of feature embeddings in ViTs leads to substantial computational overhead when performing SVD. This motivates us to explore more efficient methods for extracting orthogonal bases from feature subspaces in ViTs.
@@ 2.2 Parameter-Efficient Fine-Tuning in CL
Parameter-efficient fine-tuning (PEFT) methods enable the adaptation of pre-trained models to downstream tasks by fine-tuning only a small number of trainable parameters. Among these, prompt-tuning [10] has demonstrated strong robustness in learning sequential tasks and has achieved remarkable success in continual learning benchmarks, significantly outperforming traditional continual learning schemes. The pioneering work L2P [26] addresses continual learning challenges by introducing a prompt pool that matches the top-$k$ relevant queries, serving as supplementary inputs to facilitate feature alignment in the pre-trained model. The follow-up study, DualPrompt [25], takes a step further by employing task-invariant G-Prompts and task-specific E-Prompts to capture both shared and task-specific knowledge across different tasks. S-Prompt [24] independently learns prompts across domains, using K-means clustering on features of the training data and a K-NN algorithm to match test data with domain-specific prompts. CodaPrompt [20] shifts the focus from fixed instance-specific prompts to a set of prompt components, selecting a weighted combination of these components for each data sample. Moreover, a series of studies [16, 5, 28] fine-tune expandable prompt adapters to enhance model learnability. Although prompt-tuning methods have achieved remarkable success in continual learning, the aforementioned schemes all require forwarding training and testing data to the query function (typically the original pre-trained model) for extracting features before fine-tuning and inference, which causes longer fine-tuning and inference times.
Recently, PGP [17] and VPT-NSP [15] adapt the idea of GPM [19] in prompt-tuning to mitigate the forgetting phenomenon by implementing orthogonal projection on the gradients of the prompt pool, but inheriting all drawbacks in prompt-tuning. Similarly, InfLoRA [12] marks the first attempt to apply gradient projection in low-rank adaptation for ViT models, storing gradient directions from learned tasks during fine-tuning. However, InfLoRA introduces significant computational overhead due to performing singular value decomposition (SVD) on the high-dimensional gradient subspace, which is computationally more expensive than the original LoRA. Moreover, the orthogonal bases derived from the gradient subspace cannot be utilized to infer task identities during inference, as gradients are not available at that stage. This challenge motivates us to explore the use of gradient projection based on the orthogonal bases in the feature subspace, enabling the modulation of dynamic memory during inference for improving average test performance.
@@ 3.1 Continual Learning Problem Setting
Given a pre-trained model with backbone parameters $\mathbf{W}_{0}$, we aim to fine-tune the model by adding adapters $\mathcal{A}_{l}$ in the $l$-th layer of the model and the classifier $\mathcal{F}$ to fit a sequence of domain data $\mathcal{D}_{t}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{|\mathcal{D}_{t}|}$, where $\mathbf{x}_{i}$ denotes data samples and $\mathbf{y}_{i}\in\mathcal{Y}_{t}$ denotes corresponding labels in the $t$-th task. In the challenging class-incremental setting, there is no intersection between the label sets from two different tasks as $\mathcal{Y}_{t_{1}}\cap\mathcal{Y}_{t_{2}}=\emptyset$ for all $t_{1}\not=t_{2}$. When learning a new task, access to the old task data might become unavailable due to storage constraints. The objective function in continual learning is to minimize the empirical risk of the unified adapters $\mathcal{A}_{1:L}$ integrated in $L$ layers of the pre-trained model that performs inference on the sequence of data from $T$ various tasks as follows:
@@ 3.2 Multi-Head Attention Block
Vision transformer (ViT) models break down images into $n$ patches and flatten these patches into patch embeddings with dimension $d$. The encoder of a ViT consists of a sequence of multi-head attention (MHA) blocks containing key, query and value weights $\mathbf{W}^{q},\mathbf{W}^{k}$ and $\mathbf{W}^{v}$ for mapping input activation signals $\mathbf{a}^{(l)}$ into $\mathbf{Q}^{(l)}$, $\mathbf{K}^{(l)}$ and $\mathbf{V}^{(l)}$ and obtaining the output signals $\mathbf{h}^{(l)}$ by computing
@@ 3.3 Low-Rank Adaptation
Low-rank adaptation (LoRA) is a parameter-efficient fine-tuning method that enables reducing memory consumption by assigning learnable low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d}$ and $\mathbf{B}\in\mathbb{R}^{d\times r}$ parallel to the frozen pre-trained weights $\mathbf{W}_{0}\in\mathbb{R}^{d\times d}$ into each layer of the model as follow,
@@ 4.1 Orthogonal Adapter
The milestone work GPM [19], which uses orthogonal gradient projection to mitigate forgetting, involves flattening feature maps extracted by convolutional kernels into vectors and performing singular value decomposition (SVD) on these vectors to obtain orthogonal feature bases. However, vectorizing the patch embedding of a ViT with dimensions $(n,d)$ necessitates extensive computation in SVD, particularly when dealing with high-resolution inputs. In vision transformers, feature embeddings are often redundant for classification tasks, as only the first embedding (commonly referred to as the class token) is passed to the classifier for prediction.
@@
where $\Gamma$ is a diagonal matrix (the derivation is deferred to Appendix A.1). To preserve the value of class-token in $\mathbf{h}^{(L)}$ output by the last layer, we must restrict the value of $\Delta\mathbf{h}^{(l)}_{1}\approx\mathbf{0}$ (the change of first row in $\mathbf{h}^{(l)}$) for each layer so the class-token of the same test sample from old tasks can be preserved after fine-tuning on the new task. Since $\mathbf{Q}^{(l)}$ is unchanged with the frozen weight $\mathbf{W}_{0}^{q}$, we need to project $\left(\Delta\mathbf{O}_{t+1}^{k}\right)^{\top}$ into the subspace orthogonal to the subspace spanned by $\mathbf{k}^{(l)}:=\mathbf{Q}_{1}^{(l)}$, denoting the first row of $\mathbf{Q}^{(l)}$. Meanwhile, $\Delta\mathbf{O}_{t+1}^{v}$ must be orthogonal to the subspace of $\mathbf{v}^{(l)}:=\mathbf{S}_{1}^{(l)}$, the first row of $\mathbf{S}^{(l)}$. Following the strategy in GPM [19], we randomly sample $m$ data points from the current task after fine-tuning on the $t$-th task and input these $m$ samples to the model for obtaining embedding matrices $\tilde{\mathbf{K}}^{(l)}\in\mathbb{R}^{m\times d}$ consisting of $\{\mathbf{k}_{i}^{(l)}\}_{i=1}^{m}$ and $\tilde{\mathbf{V}}^{(l)}\in\mathbb{R}^{m\times d}$ consisting of $\{\mathbf{v}_{i}^{(l)}\}_{i=1}^{m}$. We update the new feature matrices $\boldsymbol{\Phi}_{t}^{k}$ and $\boldsymbol{\Phi}_{t}^{v}$ by extracting the orthogonal bases of $\tilde{\mathbf{K}}^{(l)}$ and $\tilde{\mathbf{V}}^{(l)}$ using SVD and concatenating them into the previous feature matrices $\boldsymbol{\Phi}_{t-1}^{k}$ and $\boldsymbol{\Phi}_{t-1}^{v}$. With the feature matrices obtained on the $t$-th task, we can project the updates $\Delta\mathbf{O}_{t+1}^{k}$ and $\Delta\mathbf{O}_{t+1}^{v}$ by

### v4

@@ Abstract
We revisit continual learning (CL), which enables pre-trained vision transformers (ViTs) to sequentially fine-tune on new downstream tasks over time. However, as the scale of these models increases, catastrophic forgetting remains a more serious challenge. Recent studies highlight a crossover between CL techniques and parameter-efficient fine-tuning (PEFT), which focuses on fine-tuning only a small set of trainable parameters to adapt to downstream tasks, such as low-rank adaptation (LoRA). Several LoRA-based CL methods have been proposed to avoid forgetting by decoupling representations learned from previous tasks and those learned for new tasks, achieving a fixed balance between stability and plasticity, yet lacking mechanisms to dynamically adjust learned representations over time. To address this gap, we propose a novel PEFT-CL method called Dual Low-Rank Adaptation (DualLoRA), which introduces both an orthogonal LoRA adapter and a residual LoRA adapter parallel to pre-trained weights in each layer. These components are orchestrated by a dynamic memory mechanism to strike a balance between stability and plasticity. Additionally, we propose a scheme to predict task identity with confidence and calibrate the model’s outputs accordingly. On ViT-based models, we demonstrate that DualLoRA offers significant advantages in accuracy, inference speed, and computation efficiency in training over existing CL methods across multiple benchmarks.
@@ 1 Introduction
Continual learning (CL) [25], which aims to train models on a sequence of tasks, often suffers from catastrophic forgetting—a substantial degradation in performance on previously learned tasks when adapting to new ones. This challenge persists even when continually fine-tuning vision foundation models [28, 29], despite their strong generalization capability and robustness. *Corresponding Author: Lingjuan Lyu Replay-based CL methods [1] attempt to alleviate forgetting by retaining a subset of data from prior tasks as exemplars; however, such approaches are often impractical due to storage constraints and data retention policies. Alternatively, architecture-based CL methods [16] mitigate interference by allocating task-specific parameters and relying on task identifiers during training. Nevertheless, modifying the architecture of vision foundation models for downstream adaptation may undermine their pre-trained representations, as these weights are learned under a fixed architectural design. Moreover, many of these methods assume access to task identity at inference time—an assumption that is frequently violated in real-world scenarios.
@@
Figure 1: PEFT-based continual learning schemes dominate the ImageNet-R dataset in recent years.
@@
Recently, parameter-efficient fine-tuning (PEFT) techniques have attracted increasing attention for their ability to adapt foundation models to downstream tasks by updating or introducing only a small number of trainable parameters. Beyond computational efficiency, PEFT methods have also demonstrated notable robustness in mitigating catastrophic forgetting under sequential task adaptation. In particular, prompt-tuning [28, 29, 24] and low-rank adaptation (LoRA) [14] have emerged as two widely adopted PEFT paradigms, achieving strong performance in continual learning and substantially outperforming conventional CL approaches on established benchmarks. Prompt-tuning methods learn a prompt pool that matches an input image with a set of prompt vectors to align image features with patch tokens. However, existing prompt-based CL approaches rely on the original pre-trained encoder as the query function. Consequently, image tokens must be processed through the backbone network twice, resulting in considerable computational overhead and increased inference latency.
In contrast, LoRA-based methods typically require fewer trainable parameters to achieve comparable performance on domain-specific tasks and enable faster inference than prompt-tuning. Nevertheless, vanilla LoRA [9] struggles in continual learning settings due to severe interference across sequential tasks. Inspired by gradient projection techniques [23], InfLoRA [14] takes an initial step toward addressing this issue by initializing LoRA adapters for new tasks within a subspace orthogonal to the gradient subspace of previously learned tasks. Building upon this idea, a series of follow-up studies [7, 33, 21, 30, 18] move beyond the strict orthogonality constraint imposed by InfLoRA and explore diverse forms of separated learning spaces, such as almost-orthogonal subspaces, principal–residual subspaces, and shared–specific subspaces, to better balance the stability–plasticity trade-off in continual learning. However, these LoRA-based CL methods still rely on a fixed set of adapted model weights during inference and do not incorporate dynamic adaptation of learned representations, which could further mitigate catastrophic forgetting.
To address this limitation, we propose a novel continual learning framework termed dual low-rank adaptation (DualLoRA). DualLoRA integrates an orthogonal adapter and a residual adapter into each layer of pre-trained vision transformers (ViTs). Specifically, the orthogonal adapter $\mathbf{O}$ is updated exclusively along directions orthogonal to the feature subspaces extracted from previously learned tasks, while the residual adapter $\mathbf{R}$ is updated within a task-specific subspace spanned by residual bases derived from prior tasks. This design promotes stability—by preventing interference with previously acquired knowledge through orthogonal adapters—while simultaneously enhancing plasticity via residual adapters that facilitate efficient adaptation to new tasks.
DualLoRA efficiently extracts orthogonal bases from the feature subspaces of past tasks and projects the updates of $\mathbf{O}$ and $\mathbf{R}$ using matrices constructed from these bases. Moreover, these bases are further leveraged at inference time to dynamically modulate the residual adapters based on input-specific task relevance, thereby suppressing components that may degrade test performance. We refer to this inference-time mechanism as dynamic memory (DM), as illustrated in Fig. 2. Extensive experimental results demonstrate that DualLoRA consistently outperforms existing PEFT-based continual learning methods across diverse benchmarks, without incurring significant additional computational or memory overhead. The main contributions of this paper are summarized as follows:
@@ 2.1 Gradient Projection in CL
Gradient projection is widely employed in continual learning to mitigate catastrophic forgetting by updating parameters in directions that minimize interference with previously learned tasks. OGD [4] was the first to implement gradient descent in directions orthogonal to the stored gradient directions computed from previous tasks. The follow-up work GPM [23] extracts orthogonal bases of task representations from randomly selected training data via singular value decomposition (SVD). A subsequent study, TRGP [15], introduces the concept of a trust region, allowing partial reuse of selected bases from previous tasks. FSDGPM [3] further evaluates the importance of bases in GPM by assessing the sharpness of the loss landscape, assigning weights to the bases in the projection matrix according to their relative importance. Due to the computational expense of determining loss landscape sharpness, SGP [22] offers an alternative by using accumulated singular values as an importance indicator to scale the projection matrix in GPM. Additionally, several studies [26, 11, 13, 31, 2] have explored relaxing the orthogonality constraints and optimizing the relaxation factor to enhance performance. However, these approaches are primarily developed for relatively simple models such as CNNs and face significant challenges when applied to advanced architectures like ViTs. For instance, the high dimensionality of feature embeddings in ViTs leads to substantial computational overhead when performing SVD. This motivates us to explore more efficient methods for extracting orthogonal bases from feature subspaces in ViTs.
@@ 2.2 Parameter-Efficient Fine-Tuning in CL
Parameter-efficient fine-tuning (PEFT) methods adapt pre-trained models to downstream tasks by fine-tuning only a small number of parameters. Among them, prompt-tuning [12] has shown strong robustness in sequential learning and achieved remarkable success in continual learning benchmarks, significantly outperforming traditional continual learning schemes. The pioneering work L2P [29] addresses continual learning by introducing a prompt pool that selects the top-$k$ relevant queries as supplementary inputs to facilitate feature alignment in the pre-trained model. DualPrompt [28] further introduces task-invariant G-Prompts and task-specific E-Prompts to capture both shared and task-specific knowledge across tasks. S-Prompt [27] learns domain-specific prompts by applying K-means clustering to training features and using a K-NN algorithm to match test samples with corresponding prompts. CodaPrompt [24] shifts from fixed instance-specific prompts to a set of prompt components, selecting a weighted combination for each sample. Moreover, several studies [19, 6, 32] propose fine-tuning expandable prompt adapters to improve model learnability. Despite their success, these methods require forwarding training and testing samples through a query function (typically the original pre-trained model) to extract features before fine-tuning and inference, resulting in longer training and inference times.
Recently, PGP [20] and VPT-NSP [17] adapt the idea of GPM [23] to prompt-tuning by applying orthogonal projection to prompt gradients to mitigate forgetting, but inherit the drawbacks of prompt-tuning. Similarly, InfLoRA [14] is the first to apply gradient projection to low-rank adaptation for ViT models by storing gradient directions from learned tasks during fine-tuning. Several follow-up studies further extend LoRA-based continual learning. CL-LoRA [7] decomposes LoRA parameters into task-shared and task-specific components to balance stability and plasticity. SD-LoRA [30] introduces a scalable decoupled LoRA design to better separate task knowledge. Split-LoRA [21] splits the gradient space into complementary subspaces to improve stability–plasticity trade-offs. BiLoRA [33] enforces almost-orthogonal parameter spaces across tasks to reduce interference. KeepLoRA [18] adapts LoRA updates via residual gradient modulation to preserve previously learned knowledge.
@@ 3.1 Continual Learning Problem Setting
Given a pre-trained model with backbone parameters $\mathbf{W}_{0}$, we aim to fine-tune the model by adding adapters $\mathcal{A}_{l}$ in the $l$-th layer of the model and the classifier $\mathcal{F}$ to fit a sequence of domain data $\mathcal{D}_{t}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{|\mathcal{D}_{t}|}$, where $\mathbf{x}_{i}$ denotes data samples and $\mathbf{y}_{i}\in\mathcal{Y}_{t}$ denotes corresponding labels in the $t$-th task. In the challenging class-incremental setting, there is no intersection between the label sets from two different tasks as $\mathcal{Y}_{t_{1}}\cap\mathcal{Y}_{t_{2}}=\emptyset$ for all $t_{1}\neq t_{2}$. When learning a new task, access to old task data might become unavailable due to storage constraints. The objective function in continual learning is to minimize the empirical risk of the unified adapters $\mathcal{A}_{1:L}$ integrated in $L$ layers of the pre-trained model that performs inference on the sequence of data from $T$ various tasks as follows:
@@ 3.2 Multi-Head Attention Block
Vision transformer (ViT) models break down images into $n$ patches and flatten these patches into patch embeddings of dimension $d$. The encoder of a ViT consists of a sequence of multi-head attention (MHA) blocks containing key, query and value weights $\mathbf{W}^{q},\mathbf{W}^{k}$ and $\mathbf{W}^{v}$ for mapping input activation signals $\mathbf{a}^{(l)}$ into $\mathbf{Q}^{(l)}$, $\mathbf{K}^{(l)}$ and $\mathbf{V}^{(l)}$ and obtaining the output signals $\mathbf{h}^{(l)}$ by computing
@@ 3.3 Low-Rank Adaptation
Low-rank adaptation (LoRA) is a parameter-efficient fine-tuning method that enables reducing memory consumption by assigning learnable low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d}$ and $\mathbf{B}\in\mathbb{R}^{d\times r}$ parallel to the frozen pre-trained weights $\mathbf{W}_{0}\in\mathbb{R}^{d\times d}$ into each layer of the model as follows,
@@ 4.1 Orthogonal Adapter
The milestone work GPM [23], which uses orthogonal gradient projection to mitigate forgetting, involves flattening feature maps extracted by convolutional kernels into vectors and performing singular value decomposition (SVD) on these vectors to obtain orthogonal feature bases. However, vectorizing the patch embeddings of a ViT with dimensions $(n,d)$ necessitates extensive computation in SVD, particularly when dealing with high-resolution inputs. In vision transformers, feature embeddings are often redundant for classification tasks, as only the first embedding (commonly referred to as the class token) is passed to the classifier for prediction.
@@
where $\Gamma$ is a diagonal matrix (the derivation is deferred to Appendix A.1). To preserve the value of the class token in $\mathbf{h}^{(L)}$ output by the last layer, we must restrict the value of $\Delta\mathbf{h}^{(l)}_{1}\approx\mathbf{0}$ (the change of the first row in $\mathbf{h}^{(l)}$) for each layer so that the class token of the same test sample from old tasks can be preserved after fine-tuning on the new task. Since $\mathbf{Q}^{(l)}$ is unchanged with the frozen weight $\mathbf{W}_{0}^{q}$, we need to project $\Delta\mathbf{O}_{t+1}^{k}$ into the subspace orthogonal to the subspace spanned by $\mathbf{k}^{(l)}:=\mathbf{Q}_{1}^{(l)}$, denoting the first row of $\mathbf{Q}^{(l)}$. Meanwhile, $\Delta\mathbf{O}_{t+1}^{v}$ must be orthogonal to the subspace of $\mathbf{v}^{(l)}:=\mathbf{S}_{1}^{(l)}$, the first row of $\mathbf{S}^{(l)}$. Following the strategy in GPM [23], we randomly sample $m$ data points from the current task after fine-tuning on the $t$-th task and input these $m$ samples to the model for obtaining embedding matrices $\tilde{\mathbf{K}}^{(l)}\in\mathbb{R}^{m\times d}$ consisting of $\{\mathbf{k}_{i}^{(l)}\}_{i=1}^{m}$ and $\tilde{\mathbf{V}}^{(l)}\in\mathbb{R}^{m\times d}$ consisting of $\{\mathbf{v}_{i}^{(l)}\}_{i=1}^{m}$. We update the new feature matrices $\boldsymbol{\Phi}_{t}^{k}$ and $\boldsymbol{\Phi}_{t}^{v}$ by extracting the orthogonal bases of $\tilde{\mathbf{K}}^{(l)}$ and $\tilde{\mathbf{V}}^{(l)}$ using SVD and concatenating them into the previous feature matrices $\boldsymbol{\Phi}_{t-1}^{k}$ and $\boldsymbol{\Phi}_{t-1}^{v}$. With the feature matrices obtained on the $t$-th task, we can project the updates $\Delta\mathbf{O}_{t+1}^{k}$ and $\Delta\mathbf{O}_{t+1}^{v}$ by